OpenStudy (anonymous):

How many solutions does the equation -2y+(4-y)=12-5y have?

OpenStudy (apoorvk):

Well, simplify it. What do you get?

OpenStudy (anonymous):

the same thing on both sides, -5y+12=12-5y. So would it be no solution?

OpenStudy (anonymous):

-2y + 4 - y = 12 - 5y Combining and gathering y terms on one side and then evaluating gives you: -3y + 4 = 12 - 5y -3y + 5y = 12 - 4 2y = 8 y = 4 So, one solution.

OpenStudy (anonymous):

Unless, what you meant to type was -2y + 3(4-y). :)

OpenStudy (anonymous):

oh, yeah.. sorry about that! haha. X)